solve for x: log(x-2) + log(x+2) = 1
Thanks

Answer by stanbon(75887)

solve for x: log(x-2) + log(x+2) = 1
--------
log[(x-2)(x+2)] = 1
---
x^2-4 = 10^1
---
x^2-14 = 0
---
(x-sqrt(14)(x+sqrt(14) = 0
Possible solutions:
x = sqrt(14) or x = -sqrt(14)
---
Check in the original equation:
x = sqrt(14)
===================
Cheers,
stan H.